OpenStudy (anonymous):

How to simplify [(x^2+4)^(1/3)-x(1/3)(x^2+4)^(-2/3)]/(x^2+4)^(2/3)?

OpenStudy (anonymous):

\[\frac{(x^2+4)^{\frac{1}{3}}-\frac{x}{3}(x^2+4)^{-\frac{2}{3}}}{(x^2+4)^{\frac{2}{3}}}\]?

OpenStudy (anonymous):

\[\frac{(x^2+4)^{\frac{1}{3}}-x^{\frac{1}{3}}(x^2+4)^{-\frac{2}{3}}}{(x^2+4)^{\frac{2}{3}}}\]?

OpenStudy (anonymous):

yes, that's the original problem.

OpenStudy (anonymous):

whichever one, multiply top and bottom by \((x^2+4)^{\frac{2}{3}}\) i think that is all you can do

OpenStudy (anonymous):

Umm... that looks complicated..Thanks for your help!